15. Suppose 0.003 moles of KOH are added to a 1.0 L solution of 0.050 moles of NH3 and 0.050 moles of NH4. What is the pH ofshow work please!
Answer #1

Ky z lokOn loxioly oxioly Soitto = 1.786*105 pkb =-109 kb -10g (1-786x103) 34-75 E K0 H3 = LOH33 770.003 50.603M Lw #] = bois